Simultaneous equations.
A food stall owner at a football game knows that hot chips are three times as popular as meat pies and 4 times as popular as Chiko rolls. During half-time in the game, the fast-food outlet served 209 people who each bought one item. How many serves of each food item were sold during this period
1
Expert's answer
2019-06-03T13:56:44-0400

Let h be the number of hot chips, m - the number of meat pies and c - the number of Chico rolls.

"h=3m,\; h=4c\;\;and\;\;h+m+c=209."

So, "h+\frac{h}{3}+\frac{h}{4}=209".

Thus, "\frac{19h}{12}=209\;\;or\;\;h=132."

"m=\frac{132}{3}=44,\;\;c=\frac{132}{4}=33."